OpenStudy (anonymous):

what is the simplest form of 10 7/6?

OpenStudy (isaiah.feynman):

Can you make that an improper fraction first?

OpenStudy (anonymous):

yes

OpenStudy (isaiah.feynman):

Please do it. :)

OpenStudy (anonymous):

67/6

OpenStudy (isaiah.feynman):

Looks like that's the simplest form.

hero (hero):

7/6 = 1 + 1/6 so 10 and 7/6 = 10 + 1 + 1/6 = 11 + 1/6 = 11 and 1/6

OpenStudy (anonymous):

thank so much